Find the no of carbon atom in a 60 grm pencil containing 60℅graphite?
1
Expert's answer
2018-07-24T06:29:25-0400
m(carbon) =m(pencil)×C% (graphite) / 100 %

m=60×60/100

m=36 grm

So, the pencil contained 36 g of graphite. Graphite is composed of complete carbon atoms. The atomic mass of 1 mole of carbon is 12 grams. Considering this, we create the following proportion

12 grm – 6,02 × 10^23

36 grm – X

X=36 × 6,02 × 10^ 23 / 12

X= 18,06 × 10^ 23


So the answer is 18,06 × 10^ 23